# cuVSLAM: CUDA accelerated visual odometry and mapping

### v1

@@ 1 Introduction
We evaluate cuVSLAM on standard benchmarks including KITTI, EuRoC, and TUM RGB-D, demonstrating state-of-the-art accuracy while maintaining real-time performance on edge computing platforms. Our results show that cuVSLAM achieves average trajectory errors below 1% on the KITTI odometry benchmark and mean position errors under 5cm on the EuRoC dataset, while processing frames in real-time on NVIDIA Jetson platforms.
@@ 2.2.2 Multicamera
Since observations from multiple cameras contribute to pose estimation, cuVSLAM requires images to be synchronized. This is a minor restriction, as many modern cameras provide hardware synchronization capabilities, such as Intel RealSense RealSense_2018.
@@ 3.1 System Performance and Hardware Utilization
We evaluated cuVSLAM’s performance in real-time live execution to measure embedded hardware utilization, as well as using recordings from a warehouse environment to measure visual tracking execution time on embedded and workstation hardware. During these experiments, cuVSLAM processed feature-rich images captured either from a wheeled robot or from handheld camera movements, representing typical usage scenarios.
Table 1: cuVSLAM Performance Evaluation. The left side shows per-call processing times measured on RTX4090 (Desktop) and Orin AGX 64gb (Jetson), all images in datasets have 768×480 resolution (except mono-depth mode with resolution 640×480). The right side summarizes hardware utilization (CPU and GPU) measured on the Jetson platform running cuVSLAM via the Python API with live RealSense cameras at 640×480 resolution and 30 FPS
. Configuration Track call time, ms Jetson HW utilization Desktop Jetson CPU % GPU % Mono 0.9 2.7 NA NA Mono-Depth 5.9 13.9 2.4 45.0 Stereo 0.4 2.7 0.4 1.7 Stereo-Inertial 1.3 3.8 Multicamera (2-stereo) 0.8 2.7 1.1 5.9 Multicamera (3-stereo) 1.4 2.8 1.5 7.1 Multicamera (4-stereo) 2.1 2.8 NA NA
The left side of Tab. 1 summarizes execution times for the cuVSLAM visual tracking function, excluding image acquisition and result post-processing overhead. Performance was measured on two hardware configurations: a desktop workstation equipped with an Intel i7-14700 CPU and NVIDIA RTX 4090 GPU, and an embedded single-board Jetson Orin AGX 64GB platform running in MAXN power mode.
The right side of Tab. 1 presents CPU and GPU utilization percentages for cuVSLAM visual odometry executed through the Python API with RealSense D435/D455 cameras running at 30 FPS. These measurements were conducted live, in real-time. To isolate cuVSLAM’s hardware resource requirements, we performed two controlled experiments: first, we measured baseline system utilization with the RealSense camera in motion but without cuVSLAM processing input data; then, we repeated the identical camera movement pattern with cuVSLAM visual tracking enabled. Resource utilization metrics were sampled at 15 ms intervals for identical durations in both experiments. cuVSLAM’s net resource consumption was calculated by subtracting baseline measurements from total utilization and averaging the results. This approach accurately reflects the additional computational overhead introduced by the library. Note that these measurements were performed in a controlled environment, excluding other resource-intensive operations such as visualization, complex image processing, or data recording, which could otherwise skew results by overloading the system.
@@ 3.2 Benchmarking
TartanAir ORBSLAM3 - - -
DPVO 12.73 0.79 12.84
cuVSLAM Odom 0.44 1.78 0.44
cuVSLAM SLAM 0.49 1.85 0.50
@@
Multi-Stereo R2B ORBSLAM3 - - -
DPVO 5.35 0.81 12.30
cuVSLAM Odom 0.18 1.15 0.28
@@
The cuVSLAM operation modes were evaluated using several publicly available datasets: Mono-Depth mode was validated on the ICL-NUIM Handa_2014 and TUM RGB-D Sturm_2012 datasets; Stereo mode was tested on the EuroC Burri_2016, KITTI Geiger_2013, and TartanAir Wang_2020 datasets; and Stereo-Inertial mode was assessed using the EuroC dataset. Standard metrics were employed to quantify the accuracy of the estimated trajectories: Average Relative Translation and Rotation Errors proposed in Geiger_2013, as well as Absolute Pose Error with trajectory alignment and scale correction Sturm_2012 using the EVO library Grupp_2017.
To ensure accurate and representative comparisons, clearly identified outlier sequences were excluded. For instance, the EuroC V203 sequence was included only in Stereo-Inertial evaluations, and certain sequences from the TartanAir and TUM RGB-D datasets were omitted. A complete list of the sequences used from each dataset is provided in Appendix A.
To rigorously validate cuVSLAM’s multi-camera capabilities, we collected a proprietary Multi-Stereo R2B dataset from office and warehouse environments using the Nova Carter wheeled robot equipped with four stereo cameras. Example scenes are shown in Fig. 7, and trajectories are illustrated in Fig. 10. A detailed description of this dataset is presented in Appendix B. Ground truth trajectories for these evaluations were obtained using high-precision lidar localization.
The use of a proprietary dataset was necessary because cuVSLAM’s multi-camera mode requires images captured by hardware-synchronized cameras, which could be combined in multiple stereo pairs. Currently, publicly available multi-camera datasets generally lack either hardware-synchronized cameras suitable for stereo pair arrangements NCTL_2015, Hilti_2024 or employ software synchronization that does not ensure simultaneous image capture M2dgr_2021, CODa_2023.
We anticipate that recent growth in availability and adoption of synchronized multi-stereo camera systems from various vendors will encourage creation of new public multi-camera datasets, facilitating further development, testing, and implementation of multi-camera SLAM systems in robotic applications.
Figure 7: Sample camera views from the R2B dataset captured across diverse environments including large office spaces, industrial warehouses, and narrow corridors in both office and storage areas. The images show simultaneous captures from 4 stereo cameras (displaying only the left image from each stereo pair) taken by the Carter Robot
Tab. 2 presents comprehensive benchmarking results across popular datasets from various domains. For comparative analysis, we include performance metrics for the classical computer vision-based ORB-SLAM3 and the deep learning-based DPVO alongside cuVSLAM. Each library was configured appropriately for the dataset domain: ORB-SLAM3 and cuVSLAM operated in their corresponding specialized modes (monocular-depth, stereo, or inertial), while DPVO was consistently used in monocular mode across all evaluations (utilizing the front left camera for stereo and multi-camera datasets).
@@ 4 Discussion
cuVSLAM demonstrates robust performance across diverse robotic platforms, including automotive vehicles, aerial drones, wheeled robots, and handheld camera setups. Our evaluation covers both real-world and synthetic datasets, providing a comprehensive assessment of the system’s capabilities under varied operational conditions.
The presented results reflect cuVSLAM’s out-of-the-box performance without parameter fine-tuning, relying solely on standard sensor calibration parameters and stable image acquisition. This approach enables successful operation with various camera types and leaves room for potential improvements by end users. For example, in our TUM-RGBD dataset evaluation, raw depth data were used without any preprocessing, despite known limitations in depth quality for this dataset.
4.1 Hardware Scalability and Implementation
cuVSLAM exhibits excellent scalability across hardware platforms and provides effective integration through user-friendly ROS2111https://github.com/NVIDIA-ISAAC-ROS/isaac_ros_visual_slam and Python [^2] https://github.com/NVlabs/PyCuVSLAM APIs. Although this paper primarily considers scenarios involving multiple USB-connected RealSense cameras operating at 30 FPS, the measured visual tracking execution times and hardware utilization indicate significant potential for more demanding configurations. This is confirmed by practical experience, as cuVSLAM has been successfully deployed to process 8 Full HD distorted RGB images at 30 FPS from 4 stereo cameras in real-time on a Jetson Orin AGX as part of the ISAAC Perceptor framework [^3] https://developer.nvidia.com/isaac/perceptor, which simultaneously runs other resource-intensive components, such as NVBLOX Millane_2023 and Nav2 Macenski_2020.
4.2 Advantages and Research Directions
We advocate for wider adoption of multi-camera visual tracking in robotics based on two principal observed advantages: enhanced trajectory reliability in feature-poor environments in odometry mode, and increased loop closure detection rates in SLAM mode. Fig. 8 demonstrates trajectory reliability through a practical test conducted using a Carter robot equipped with four stereo cameras. In this experiment, the robot moved continuously while its cameras were randomly covered with non-transparent film for intervals of 20–60 seconds, ensuring at least one stereo pair remained uncovered at given time. Despite these challenging conditions, the odometry trajectory remained stable without jumps or discontinuities, validating the system’s robustness to partial visual obstruction.
Figure 8: Occlusion resilience test results using the Carter robot. The figure shows the continuous trajectory successfully maintained by cuVSLAM multicamera visual odometry despite random camera occlusions using non-transparent film. Numbered image rows correspond to instantaneous views from the robot’s four stereo cameras (showing only the left image from each stereo pair) at the respective numbered locations along the trajectory.
Additionally, the increased loop closure detection rate resulting from an expanded collective field of view is demonstrated in Fig. 9, comparing the number of loop closure events between single-stereo and four-stereo camera setups for the same trajectory. This higher frequency of camera pose adjustments yields a quantitative improvement of approximately 40% in SLAM accuracy metrics compared to odometry mode, as demonstrated in Tab. 2 using the Multi-Stereo R2B dataset.

### v2

@@ 1 Introduction
We evaluate cuVSLAM on standard benchmarks, demonstrating state-of-the-art accuracy while maintaining real-time performance on edge computing platforms. Our results show that cuVSLAM achieves average trajectory errors below 1% on the KITTI odometry benchmark and position errors under 5cm on the EuRoC dataset, while processing frames in real-time on NVIDIA Jetson platforms. Multi-Stereo mode significantly improves accuracy and robustness on challenging sequences compared to single stereo cameras, as validated through both simulated and real-world dataset evaluations.
@@ 2.2.2 Multi-Stereo
Since observations from multiple cameras contribute to pose estimation, cuVSLAM requires images to be synchronized. This is a minor restriction, as many modern cameras provide hardware synchronization capabilities, such as Intel RealSense Grunnet-Jepsen et al. (2018).
@@ 3.1 System Performance and Hardware Utilization
We evaluated cuVSLAM’s performance through two complementary approaches: real-time execution on embedded hardware to measure resource utilization and single-frame processing time, and offline processing of warehouse environment recordings on workstation hardware. The test scenarios involved processing feature-rich images captured from either a wheeled robot navigating a warehouse environment or handheld camera movements. To ensure accurate measurements, all experiments were conducted on systems dedicated solely to cuVSLAM execution, without concurrent resource-intensive operations such as visualization, complex image processing, or data recording that could potentially skew the results.
Table 1 summarizes the key performance metrics, with comprehensive results provided in Appendix A.2. The left portion of Table 1 presents execution times for cuVSLAM’s visual tracking function, excluding image acquisition and post-processing overhead. These measurements were obtained using recorded color images at 768×480 resolution (with the exception of mono-depth mode at 640×480 resolution) on two distinct hardware configurations: a desktop workstation featuring an Intel i7-14700 CPU with NVIDIA RTX 4090 GPU, and an embedded Jetson Orin AGX 64GB single-board computer operating in MAXN power mode.
Table 1: cuVSLAM Performance Evaluation. Left: Per-call processing times measured on RTX 4090 (Desktop) and Jetson AGX Orin (64 GB) using datasets with image resolution 768×480 (except mono-depth mode at 640×480). Right: CPU and GPU utilization on Jetson AGX Orin during live operation with RealSense cameras at 640×480 resolution and 60 FPS (except stereo-inertial mode at 30 FPS).
. Configuration Track call time, ms Jetson HW utilization Desktop Jetson CPU % GPU % Mono 0.9 2.7 NA NA Mono-Depth 5.9 15.1 2.6 55.0 Stereo-Inertial 1.3 3.8 1.3 2.2 Stereo 0.4 1.8 5.5 1.7 Multicamera (2-stereo) 0.8 2.0 8.3 9.0 Multicamera (3-stereo) 1.4 2.3 8.3 11.0 Multicamera (4-stereo) 2.1 NA NA NA
The right portion of Table 1 details CPU and GPU utilization percentages for cuVSLAM visual odometry executed through ISAAC ROS 3.2 with RealSense D435/D455 stereo cameras. All live tests utilized cameras configured at 640×480 resolution and 60 FPS, except for Stereo-Inertial mode, which ran at 30 FPS with IMU data at 200 Hz. The reduced framerate was necessary to ensure valid IMU-based integration, providing 6-7 IMU measurements between each stereo image pair. To accurately isolate cuVSLAM’s hardware requirements, we employed a controlled experimental methodology: we first established baseline system utilization with the RealSense camera in motion but without cuVSLAM processing, then repeated the similar camera movement pattern with cuVSLAM visual tracking active. Resource utilization metrics were sampled at 15 ms intervals for equal durations in both experiments. The net resource consumption attributable to cuVSLAM was calculated by subtracting baseline measurements from total utilization values and averaging the results. This methodology provides an accurate assessment of the computational overhead introduced by the library. Additional evaluation details, including performance benchmarks for cuVSLAM executed via Python API on entry-level Jetson Orin Nano hardware, are presented in Appendix A.2.
@@ 3.2 Benchmarking
We evaluated cuVSLAM’s different operation modes using several publicly available datasets. The Mono-Depth mode was validated on the ICL-NUIM Handa et al. (2014) and TUM RGB-D Sturm et al. (2012) datasets. The Stereo mode was tested on the EuRoC Burri et al. (2016) and KITTI Geiger et al. (2013) datasets. The Stereo-Inertial mode was assessed using the EuRoC dataset. The Multi-Stereo mode was evaluated on simulated TartanAir V2 Wang et al. (2023) and TartanGround Patel et al. (2025) datasets and our proprietary R2B real-world dataset.
To quantify trajectory estimation accuracy, we employed standard metrics: Average Relative Translation and Rotation Errors Geiger et al. (2013), and Absolute Pose Error with trajectory alignment and scale correction Sturm et al. (2012) using the EVO library Grupp (2017).
@@
Table 2 presents comprehensive benchmarking results across popular datasets from various domains. For comparative analysis, we include performance metrics for the classical computer vision-based ORB-SLAM3 and the deep learning-based DPVO alongside cuVSLAM. Each library was configured according to the dataset characteristics: ORB-SLAM3 and cuVSLAM operated in their corresponding specialized modes (monocular-depth, stereo, or stereo-inertial), while DPVO consistently used monocular mode across all evaluations, utilizing the front-left camera for stereo datasets. Since scale corrections were applied to final trajectories rather than successive pose estimations, translation results may not be meaningful for some datasets; therefore, relative rotation error should be considered the primary metric.
To ensure accurate and representative comparisons, we excluded clearly identified outlier and corrupted sequences. For instance, the EuRoC V203 sequence was included only in Stereo-Inertial evaluations, and only a subset of TUM RGB-D sequences were used (see Table 4). For TartanGround datasets, not all sequences provide the full complement of cameras or are available for download. Additionally, many sequences lack overlapping trajectories, preventing loop closure detection—a critical component for evaluating SLAM performance versus odometry. Consequently, we validated SLAM mode on selected subsets of TartanGround (see Table 7) and TartanAir V2 (see Tables 9 and 9) datasets.
Figure 7: Sample camera views from the R2B dataset captured across diverse environments including large office spaces, industrial warehouses, and narrow corridors in both office and storage areas. The images show simultaneous captures from 4 stereo cameras (displaying only the left image from each stereo pair).
While the new Tartan datasets provide free camera and quadruped robot movements in diverse simulated environments, we needed to rigorously validate cuVSLAM’s multi-camera capabilities in real-world scenarios. We therefore collected a proprietary Multi-Stereo R2B dataset in office and warehouse environments using the Nova Carter wheeled robot equipped with four stereo cameras. The example views from the R2B dataset are shown in Fig. 7. This was our only available option for validating this cuVSLAM mode on real data, as publicly available real-world multi-camera datasets either lack hardware-synchronized cameras suitable for stereo pair arrangements Carlevaris-Bianco et al. (2015); Nair et al. (2024) or employ software synchronization that cannot ensure simultaneous image capture Yin et al. (2021); Zhang et al. (2023).
Table 3 presents cuVSLAM’s results in Multi-Stereo Camera Mode. Notably, results achieved on the real R2B dataset closely match those from similar domains in the TartanGround dataset, confirming the validity of both datasets and results. For TartanAir V2, we performed detailed validation only on the Hard dataset, which presents greater challenges due to fast six-degrees-of-freedom motion and higher translational and angular velocities. Results for the Easy mode are comparable to those for TartanGround, with a detailed comparison available in Table 6. While TartanAir V2 results indicate room for improvement, they also demonstrate the importance of multi-camera mode for robust tracking in most environments, with some showing 2–4× improvement in multi-camera SLAM mode.
Table 3: Evaluation results for cuVSLAM Multi-Stereo Visual Odometry and SLAM
Dataset Method ATE, % ARE, deg APE (RMSE)
R2B cuVSLAM Odom 0.18 1.15 0.28
@@
TartanGround cuVSLAM Odom* 0.54 0.95 0.14
cuVSLAM Odom 0.21 0.48 0.09
cuVSLAM SLAM 0.17 0.37 0.07
TartanAir V2 (Hard) cuVSLAM Odom* 2.56 13.97 5.20
cuVSLAM Odom 2.44 13.98 5.24
cuVSLAM SLAM 2.26 12.76 4.99
We anticipate that the recent growth in availability and adoption of synchronized multi-stereo camera systems from various vendors will encourage the creation of new public multi-camera datasets, facilitating further development, testing, and implementation of multi-camera SLAM systems in robotic applications.
@@ 4 Discussion
We advocate for the wider adoption of multi-camera visual tracking in robotics based on two principal advantages observed in our experiments: enhanced trajectory reliability in feature-poor environments when operating in odometry mode, and increased loop closure detection rates in SLAM mode.
4.1 Trajectory Reliability Under Visual Occlusion
Figure 8 demonstrates the system’s trajectory reliability through a practical robustness test. We equipped a Carter robot with 4 stereo cameras and subjected it to challenging visual conditions during continuous motion. Cameras were randomly occluded with opaque film for intervals of 20–60 seconds, with the constraint that at least one stereo pair remained unobstructed at a given moment. Despite these severe visual constraints, the odometry trajectory maintained stability without exhibiting jumps or discontinuities, validating the system’s robustness to partial visual obstruction. This capability represents a significant advancement for visual tracking systems. While single-camera configurations can maintain tracking in feature-poor environments only briefly—typically relying on IMU or depth sensor fusion—our multi-camera approach preserves stable and reliable visual tracking over extended periods even with degraded visual input.
Figure 8: Occlusion resilience test results using the Carter robot. The figure shows the continuous trajectory successfully maintained by cuVSLAM multicamera visual odometry despite random camera occlusions using non-transparent film. Numbered image rows correspond to instantaneous views from the robot’s 4 stereo cameras (showing only the left image from each stereo pair) at the numbered locations along the trajectory.
4.2 Enhanced Loop Closure Detection
The expanded collective field of view afforded by multiple cameras directly translates to increased loop closure detection rates. Figure 9 presents a comparative analysis of loop closure events between single-stereo and four-stereo camera configurations along identical trajectories. The multi-camera setup demonstrates a substantially higher frequency of successful loop closures, enabling more frequent pose graph optimization. This increased loop closure rate yields quantitative improvements in localization accuracy. As shown in Table 3, evaluation on the Multi-Stereo R2B dataset reveals approximately 40% improvement in SLAM accuracy metrics compared to pure odometry mode, underscoring the practical benefits of the multi-camera approach for robust localization.
@@ 4.2 Enhanced Loop Closure Detection
5 Conclusions
This technical report demonstrates the robust real-time performance of the cuVSLAM visual tracking library across diverse robotic platforms, including automotive vehicles, aerial drones, wheeled robots, and handheld camera systems. Our comprehensive evaluation encompasses both real-world deployments and synthetic datasets, validating the system’s capabilities under varied operational conditions.
The cuVSLAM library offers several key advantages that position it as a production-ready solution for modern robotics applications:
  • Accessibility and Integration: The library is readily available through both ISAAC ROS [^1] https://github.com/NVIDIA-ISAAC-ROS/isaac_ros_visual_slam and Python [^2] https://github.com/NVlabs/PyCuVSLAM APIs, supporting deployment on both desktop and embedded NVIDIA platforms. This dual-API approach ensures flexibility for developers working across different software ecosystems.
  • Computational Efficiency: The system maintains low computational overhead while delivering high-accuracy localization, enabling seamless integration into complex robotic stacks such as NVIDIA Isaac Perceptor [^3] https://developer.nvidia.com/isaac/perceptor, which incorporates NVBLOX Millane et al. (2023) for 3D reconstruction and Nav2 Macenski et al. (2020) for autonomous navigation.
  • Minimal Configuration Overhead: Unlike many SLAM systems that require extensive parameter tuning for different scenarios, cuVSLAM provides robust out-of-the-box performance across diverse environments and sensor configurations, significantly reducing deployment time and complexity.
Looking forward, the demonstrated multi-camera capabilities and robust performance characteristics of cuVSLAM establish a strong foundation for next-generation autonomous systems that demand reliable visual localization in challenging real-world environments. The library’s proven versatility across platforms and scenarios, combined with its efficient GPU-accelerated implementation, makes it a compelling choice for both research prototypes and production robotic systems.